OpenStudy (anonymous):

What is 4b to the 4th power when b=3?

OpenStudy (wade123):

4b^4 like this right?

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ (4b)^4=(4 \times 3)^4=12^4=20,736}\)

OpenStudy (wade123):

oh ok

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

either \(\normalsize\color{black}{ (4b)^4=(4 \times 3)^4=12^4=20,736}\) or \(\normalsize\color{black}{ 4(b)^4=4 (3)^4=4 \times 81=324}\)

OpenStudy (anonymous):

Do you know which one?

OpenStudy (solomonzelman):

I am thinking when they word the problem as 4b to the forth power, it is the first one, because if it was otherwise, then they would have said " 4 times b to 4th power"

OpenStudy (anonymous):

ok thanks

OpenStudy (solomonzelman):

Anytime !

OpenStudy (triciaal):

I would use (4b)^4

OpenStudy (solomonzelman):

Nice van, bwt :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

What is 4a to the 2nd power - 5b to the 2nd power when a=5 and b=3

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ (4a)^2-(5b)^2}\) \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ (4a-5b)(4a+5b)}\) \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ (4(5)-5(3))(4(5)+5(3))}\) \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ (20-15)(20+15)}\) \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ (5)(35)}\) \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ 175}\) \(\LARGE\color{white}{ \rm │ }\)